OpenStudy (alecwilson):

Using complete sentences, write an indirect proof proving that if x = 40, then 5/6 x - 3 ≠ 22.

OpenStudy (alecwilson):

@jim_thompson5910 if you aren't busy, i am clueless

jimthompson5910 (jim_thompson5910):

Because it's an indirect proof, you have to do a proof by contradiction

jimthompson5910 (jim_thompson5910):

assume the opposite is true assume that (5/6)x - 3 = 22 is true solve for x and tell me what you get

OpenStudy (alecwilson):

30=25 which isnt true?

OpenStudy (alecwilson):

(5/6)x=25?

jimthompson5910 (jim_thompson5910):

what comes next after (5/6)x=25

OpenStudy (alecwilson):

5/6 of 25 is 83.3%?

jimthompson5910 (jim_thompson5910):

multiply both sides by the reciprocal of 5/6

OpenStudy (alecwilson):

30/30 And 30?

OpenStudy (alecwilson):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):